# UniHENN: Designing More Versatile Homomorphic Encryption-based CNNs without im2col

### v1

@@ IV-B Construction of the Convolutional Layer
The im2col encoding is a highly efficient algorithm for converting multidimensional data into matrix form to facilitate convolution operations. However, it requires precise arrangement of input matrix elements, which is challenging to achieve on ciphertexts. Therefore, TenSEAL [7], a sophisticated open-source HE-enabled ML SDK, performs this operation in a preprocessing step before encryption, allowing only a single convolution.
To overcome the limitations of im2col, we introduce a new method for computing convolutional layers that does not rely on this encoding. Our approach uses flattened input data, eliminating the need for a specific arrangement of input matrix elements and thereby increasing flexibility.
In convolutional layers, a kernel traverses an input image with each movement termed a stride. The kernel executes a convolution operation at each stride, multiplying its values with corresponding input image values and summing the results. The input image dimensions are denoted as $(W_{img},H_{img})$, and the stride as $S$. And for simplicity, we define the width and height are equal and denoted as $K$. The total number of kernel movements is given by Equation 1:
$\left(\frac{W_{img}-S}{K}\right)\left(\frac{H_{img}-S}{K}\right)$ (1)
The total number of operations executed by a convolutional layer is influenced by both the size of the input image and the kernel, as indicated by Equation 1. However, given that the input image is usually significantly larger than the kernel, we can approximate that the total number of operations primarily depends on the kernel size.
@@
As shown in Figure 4, the input data dimension for a basic convolution is two-dimensional. However, input data must be restricted to one-dimensional form in the context of HE. The input data is flattened, as demonstrated in Figure 5, using the method outlined in Section III-A. For simplicity, we assume that both the input data and the kernel are two-dimensional squares and disregard bias operations.
Figure 5: Convolution operations on the flattened input data.
In Figure 5, the number of rotation operations equals the number of elements in the kernel. The input data is incrementally increased by the original kernel size $K$, followed by a leftward rotation operation after each increment. After adding $K\times W_{img}$ cipertexts, another leftward rotation is performed, incrementing by one at each step. The detailed algorithm of these mechanism is described in Algorithm 1. All rotated ciphertexts are then multiplied by the kernel vector, which is specially structured to contain a single kernel element with all other positions set to zero. This eliminates unrelated elements in the ciphertext during convolution, as depicted in Figure 6. For 1D convolution, the logic remains the same, except that the one-dimensional data consists of a single row. The output is an array spaced by the stride interval, as shown in Figure 7.
Figure 7: Output of convolution operations on one-dimensional data.
This approach has the advantage of decoupling the number of operations from the size of the image. Instead, it performs convolutional operations on transformed two-dimensional data and limits rotations to the square of the kernel size $K$.
Previous approaches have used im2col encoding [7] or performed operations on individual data values [16]. However, the im2col method restricts the model to a single convolutional layer and requires model-specific encoding. Methods that operate on each data value are considerably slower.
@@
Procedure: *(All ciphertexts and operation are in $\mathcal{R}_{q}^{2}$)
@@
     for $p=0$ to $K-1$ do
         for $q=0$ to $K-1$ do
              $C_{rot(i,p,q)}$
              $\leftarrow Rot\left(C_{in(i)},I_{in}\times(q+W_{img}\times p)\right)$
@@
     $C_{out(o)}$
     $\leftarrow\sum_{i=0}^{CH_{in}-1}\sum_{p=0}^{K-1}\sum_{q=0}^{K-1}Mul\left(C_{rot(i,p,q)},K_{(o,i,p,q)}\right)$
@@
The number of multiplications in the convolution layer of UniHENN is $\mathcal{O}(CH_{in}\cdot CH_{out}\cdot K^{2})$, and the number of rotations in the convolution layer of UniHENN is $\mathcal{O}(CH_{in}\cdot K^{2})$. As stated in  [32], the complexity of a single multiplication is $\mathcal{O}(N\cdot D)$, and the complexity of a single rotation is $\mathcal{O}(N\cdot\log N\cdot D^{2})$. Therefore, the complexity of the convolutional layer of UniHENN is $\mathcal{O}(N\cdot D\cdot CH_{in}\cdot K^{2}\cdot(CH_{out}+\log N\cdot D))$.
@@ IV-C Construction of the Average Pooling Layer
CNNs use pooling layers to reduce the input size of convolutional layers. The most commonly used pooling layers are max pooling, min pooling, and average pooling. Max and min pooling require many multiplications [34], while average pooling requires one. Our architecture aims to provide a CNN inference with high speed, we only consider the CNN models without bootstrapping. But, the max pooling and min pooling consume lots of multiplications so it is needed to use the bootstrapping with these pooling layers. Therefore, in UniHENN, we adopt average pooling instead of max pooling and min pooling.
Originally, the average operation is a multiplication of the constant $1/c^{2}$, where $c$ is the size of the kernel. However, in UniHENN, we modify the average pooling layer to eliminate the need for constant multiplication. Our main idea is to apply the constant multiplication in one of the convolutional layers or flatten layers instead of the average pooling layer. It provides the same functionality as the general average pooling but reduces multiplication only changing the order of multiplication of $1/c^{2}$ in the previous layer. So, it can choose lower depth, resulting in efficiency in the overall operation.
The choice is determined by the layer that follows average pooling. If an activation layer follows, the constant multiplication is passed to either the flatten layer or the next convolutional layer, as this layer is linear. Using the linearity $h(x)=Ax+b$, we find $h(x/c^{2})=A(x/c^{2})+b=(1/c^{2})Ax+b$, allowing us to use the weight $(1/c^{2})A$ instead of $A$.
@@ IV-D Construction of the Fully Connected (FC) Layer
In Figure 9, the overall process of vector multiplication in UniHENN is illustrated. First, as shown in Figure 9 (a), pad zeros until the number of rows is a multiple of the output size, and then rotate the columns of the weight matrix upward, incrementing one column at a time as shown in Figure 9 (b). Next, append the segments of each array with a negative index to the end of the array. Finally, as depicted in Figure 9 (c), divide the matrix lengthwise by the output size and concatenate the parts horizontally; the weight matrix $M_{w}$ then becomes $M_{rot}$ as shown in Figure 9 (d). Note that generating an optimized weight matrix (from Figure 9 (a) to Figure 9 (d)) involves modifying the weights in plaintext. Therefore, the operation time is negligible. Algorithm 2 provides the details of the process for constructing the fully connected layer.
@@
     $V_{1}\leftarrow M_{rot(o)}[:DAT_{out}-o]+[0]\times o$
     $C_{out(o)}\leftarrow Mul(Rot(C_{in},o),V_{1})$
     $V_{2}\leftarrow[0]\times(DAT_{out}-o)+M_{rot(o)}[DAT_{out}-o:]$
     $V_{3}\leftarrow Mul(Rot(C_{in},o-DAT_{out}),V_{2})$
     $C_{out(o)}\leftarrow Add(C_{out(o)},V_{3})$
@@
for $i=1$ to $\lceil DAT_{in}/DAT_{out}\rceil-1$ do
     $C_{out}\leftarrow Add(C_{out},Rot(C_{out},i\times DAT_{in}))$
@@
$C_{out}\leftarrow Add(C_{out},B)$
return $C_{out}$
The number of multiplications in the fully connected (FC) layer of UniHENN is $\mathcal{O}(DAT_{out})$, and the number of rotations in the FC layer of UniHENN is $\mathcal{O}(DAT_{in}/DAT_{out}+DAT_{out})$. This implies that if $DAT_{in}$ is smaller than the square of $DAT_{out}$, the rotations can be denoted as $\mathcal{O}(DAT_{out})$. In this scenario, the time complexity of the FC layer depends solely on the output size of the FC layer. Thus, the complexity of the FC layer of UniHENN is $\mathcal{O}((DAT_{in}/DAT_{out}+DAT_{out})\cdot N\cdot\log N\cdot D^{2})$.
@@ IV-E1 Removing the row interval
After passing through the convolution layer or fully connected layer, the valid data in the output ciphertext will have offset, which we call row interval. We propose an algorithm to remove the row interval for two situations: the preceding layer is an average pooling layer or the preceding layer is not an average pooling. As detailed in Section IV-C, the square layer does not handle constant multiplication, so it is not considered when identifying the previous layer.
As illustrated in Figure 10, if the previous layer is an average pooling layer, the given input ciphertexts have invalid values in the intervals between each set of values. Furthermore, since we deferred the constant multiplication ($1/c^{2}$) in the pooling layer, we need to multiply $1/c^{2}$ to the ciphertexts. Section IV-E2 describes the three steps to address this issue.
If the preceding layer is either a convolutional layer or an approximate ReLU layer, it will contain some zero values between each set of non-zero values. We can exploit this to efficiently execute the flatten operation by stacking the non-zero values using rotation and addition.
@@ IV-E4 Packing multiple ciphertexts
The number of multiplications in the flatten layer of UniHENN is $\mathcal{O}(W_{in}+H_{in})$, and the number of rotations in the flatten layer of UniHENN is $\mathcal{O}(CH_{in}\cdot S_{total}+W_{in}+H_{in})$.
Therefore the complexity of flatten layer of UniHENN is $\mathcal{O}((CH_{in}\cdot S_{total}+W_{in}+H_{in})\cdot N\cdot\log N\cdot D^{2})$
@@ V-A Convolutional Layer
Let $W_{img}\times H_{img}$ be the image size and $L$ the number of total convolutional layers. Denote $(W_{in(i)},H_{in(i)})$ and $(W_{out(i)},H_{out(i)})$ as the input and output sizes of the $i$-th convolutional layer, and the kernel size is $(KW_{i},KH_{i})$, stride size is $(SW_{i},SH_{i})$, and padding size is $(PW_{i},PH_{i})$ for all $i\in\left[1,L\right]$. Additionally, let
$SH_{(i)}=\prod\limits_{j=1}^{i}SH_{j}$ (3)
Then, $H_{img}\geq H_{out(i)}\times SH_{(i)}$, where $KH_{i}\geq SH_{i}\geq 1$ for all $i\in\left[1,L\right]$.
@@
$H_{out(i)}=\left[\frac{H_{in(i)}-KH_{i}}{SH_{i}}\right]+1$ (4)
@@
$H_{out(i)}\times SH_{i}$ $=\left(\left[\frac{H_{in(i)}-KH_{i}}{SH_{i}}\right]+1\right)\times SH_{i}$ (5)
$=\left[\frac{H_{in(i)}-KH_{i}+SH_{i}}{SH_{i}}\right]\times SH_{i}$ (6)
$\leq\left[\frac{H_{in(i)}}{SH_{i}}\right]\times SH_{i}\leq H_{in(i)}=H_{out(i-1)}$ (7)
@@
Note that after passing through the $i$-th convolutional layer, the interval in each column of data is $W_{img}\times(SH_{(i)}-1)$ (detailed in  IV-B). Given $W_{img}$ columns of data or fewer, the image size of $H_{img}\times W_{img}$ is not exceeded by Theorem 1.
@@
$SW_{(i)}=\prod\limits_{j=1}^{i}SW_{j}$ (8)
Then, $W_{img}\geq W_{out(i)}\times SW_{(i)}$, where $KW_{i}\geq SW_{i}\geq 1$ for all $i\in\left[1,L\right]$.
@@ V-B Average Pooling Layer
Assuming that the kernel size in the average pooling layer is $c$, each channel uses a filter with a kernel size of $(c,c)$ and a stride of $(c,c)$. The result is then multiplied by $1/c^{2}$. We omit the multiplication of a vector consisting of 0s and 1s, allowing some non-zero but unused values (i.e., invalid values) to overlap with other values. According to Theorem 1, the number of slots in the ciphertext data remains unchanged when passing through UniHENN’s average pooling. However, to prevent invalid values from occupying the positions of other used data, an additional space of $(W_{img}+1)\times(c-1)$ is required (detailed in Section IV-C).
@@ VI-A Experimental Settings
We used a server service on the NAVER Cloud platform [42] with the following specifications: 16 vCPU cores (Intel(R) Xeon(R) Gold 5220 CPU @ 2.20GHz), 64GB of memory, and a 50GB SSD.
@@ VI-B Datasets
We utilize four different datasets for each model: MNIST [15], CIFAR-10 [26], USPS [19], and ECG [38].
@@ VI-B1 MNIST Dataset
MNIST is a dataset [15] consisting of 70,000 grayscale images of handwritten digits, each with a size of $28\times 28$ pixels. The dataset contains 10 classes, representing the digits from 0 to 9. The dataset consists of two parts: 60,000 data for training and 10,000 data for testing. We use these settings in our experiments.
@@ VI-B2 CIFAR-10 Dataset
CIFAR-10 is a dataset [26] comprising 60,000 RGB images, each with dimensions $3\times 32\times 32$ pixels. Each image is labeled with one of 10 classes. The dataset consists of two parts: 60,000 data for training and 10,000 data for testing. We use these settings in our experiments.

### v3

@@ IV-B Construction of the Convolutional Layer
The im2col encoding is a popular and efficient algorithm for transforming multidimensional data into matrix form to facilitate convolution operations. However, it requires precise arrangement of input matrix elements, which is challenging to achieve on ciphertexts. Therefore, TenSEAL [7], a sophisticated open-source HE-enabled ML SDK, performs this operation in a preprocessing step before encryption, allowing only a single convolutional layer.
To simplify this process and improve flexibility, we propose a new method for constructing convolutional layers without relying on im2col. Our approach uses flattened input data, eliminating the need for a specific arrangement of input matrix elements as follows:
  1. 1.
    Kernel and Stride: In convolutional layers, a small matrix called a kernel moves over the input image. Each movement of the kernel is called a stride. At each position, the kernel multiplies its values with the corresponding values in the input image and sums the results.
  2. 2.
    Input and Kernel Dimensions: The dimensions of the input image are denoted as $(W_{img},H_{img})$, with a stride of $S$. For simplicity, we assume the kernel has equal width and height, both represented by $K$.
  3. 3.
    Flattening the Input: In the context of HE, we flatten the two-dimensional input data into a one-dimensional form. This simplifies the convolution process under encryption.
  4. 4.
    Rotation and Multiplication: The convolution operation involves rotating the input data and multiplying it with the kernel. The number of rotations equals the number of elements in the kernel. Each element in the input data is multiplied by the corresponding element in the kernel, and the results are summed up.
  5. 5.
    Output: The output of the convolution is an array with values spaced according to the stride interval.
Figures 4 and 5 illustrate the convolution operations on two-dimensional and flattened input data, respectively.
@@
Figure 5: Convolution operations on the flattened data.
Figure 5 illustrates the convolution process: the input data is incrementally shifted by $S$, $K$ times, with leftward rotations after each shift. These rotated ciphertexts are then multiplied by a specially structured kernel vector containing a single kernel element, effectively eliminating unrelated elements during convolution. Algorithm 1 provides a detailed description of this convolution layer construction.
@@
Procedure: *(All ciphertexts and operations are in $\mathcal{R}_{q}^{2}$)
@@
     for $j=0$ to $K-1$ do
         for $k=0$ to $K-1$ do
              $C_{rot(i,j,k)}$
              $\leftarrow Rot\left(C_{in(i)},I_{in}\times(k+W_{img}\times j)\right)$
@@
     $C_{out(t)}$
     $\leftarrow\sum_{i=0}^{CH_{in}-1}\sum_{j=0}^{K-1}\sum_{k=0}^{K-1}Mul\left(C_{rot(i,j,k)},K_{(o,i,j,k)}\right)$
@@
This approach improves the convolution process by decoupling the number of operations from the image size and limiting rotations to the square of the kernel size $K$.
The number of plaintext multiplications in the convolution layer of our method is $\mathcal{O}(CH_{in}\cdot CH_{out}\cdot K^{2})$ and the number of rotations is $\mathcal{O}(CH_{in}\cdot K^{2})$. Considering the complexity of plaintext multiplication $\mathcal{O}(N\cdot L)$ and rotation $\mathcal{O}(N\cdot\log N\cdot L^{2})$ as stated in [33], the overall complexity of the convolutional layer is $\mathcal{O}(N\cdot L\cdot CH_{in}\cdot K^{2}\cdot(CH_{out}+\log N\cdot L))$.
@@ IV-C Construction of the Average Pooling Layer
CNNs use pooling layers to reduce input size, with max, min, and average pooling being common. Our architecture prioritizes high-speed CNN inference without bootstrapping. Ablation studies on a LeNet-5-like model revealed comparable accuracies for min (0.989), max (0.992), and average (0.985) pooling. In HE, max and min pooling require numerous multiplications [35], necessitating costly bootstrapping, while average pooling needs only one multiplication. Given the negligible performance difference and significant computational savings, we implemented average pooling in UniHENN.
We optimize the average pooling layer by eliminating the constant multiplication by $1/c^{2}$ (where $c$ is the kernel size). Instead, we apply this multiplication in a preceding convolutional or flatten layer. This approach maintains functionality while reducing the number of multiplications and lowering depth, enhancing overall efficiency.
The implementation depends on the subsequent layer. If followed by an activation layer, the constant multiplication is moved to the preceding flatten or convolutional layer. For a linear layer $h(x)=Ax+b$, we use $h(x/c^{2})=(1/c^{2})Ax+b$, incorporating the scaling factor into the weight matrix. This strategy preserves the average pooling effect while optimizing computational resources.
@@ IV-D Construction of the Fully Connected (FC) Layer
In Figure 8, the overall process of vector multiplication in UniHENN is illustrated. First, as shown in Figure 8 (a), pad zeros until the number of rows is a multiple of the output size, and then rotate the columns of the weight matrix upward, incrementing one column at a time as shown in Figure 8 (b). Next, append the segments of each array with a negative index to the end of the array. Finally, as depicted in Figure 8 (c), divide the matrix lengthwise by the output size and concatenate the parts horizontally; the weight matrix $M_{w}$ then becomes as shown in Figure 8 (d). Note that generating an optimized weight matrix (from Figure 8 (a) to Figure 8 (d)) involves modifying the weights in plaintext; however, the operation time is negligible.
@@
Procedure: *(All ciphertexts and operations are in $\mathcal{R}_{q}^{2}$)
$M_{rot}\leftarrow$ Apply $M_{w}$ from step 1 to 4 in Figure 8
$W_{vec}=\lceil DAT_{in}/DAT_{out}\rceil\times DAT_{out}$
$C_{sum}\leftarrow C(0)$
@@
     $V_{1}\leftarrow M_{rot(o)}[:W_{vec}-o]+[0]\times o$
     $C_{sum}\leftarrow Add(C_{sum},Mul(Rot(C_{in},o),V_{1}))$
     $V_{2}\leftarrow[0]\times(W_{vec}-o)+M_{rot(o)}[W_{vec}-o:]$
     $C_{sum}\leftarrow Add(C_{sum},Mul(Rot(C_{in},o-W_{vec}),V_{2}))$
@@
$C_{out}\leftarrow C(0)$
for $i=0$ to $\lceil DAT_{in}/DAT_{out}\rceil-1$ do
     $C_{out}\leftarrow Add(C_{out},Rot(C_{sum},i\times DAT_{out}))$
@@
$C_{out}\leftarrow Add(C_{out},B)$ return $C_{out}$
The number of plaintext multiplications of the FC layer in UniHENN is $\mathcal{O}(DAT_{out})$, and the number of rotations is $\mathcal{O}(DAT_{in}/DAT_{out}+DAT_{out})$. This means that if $DAT_{in}$ is smaller than the square of $DAT_{out}$, it can be denoted as $\mathcal{O}(DAT_{out})$. In this case, the time complexity of the FC layer depends only on the output size. Therefore, the complexity of the FC layer in UniHENN is $\mathcal{O}((DAT_{in}/DAT_{out}+DAT_{out})\cdot N\cdot\log N\cdot L^{2})$.
@@ IV-E1 Removing the Row Interval
After passing through the convolution layer or fully connected layer, the valid data in the output ciphertext will be offset, which we call row interval. We propose an algorithm to remove the row interval for two situations: whether the preceding layer is an average pooling layer or not. As detailed in Section IV-C, the square layer does not handle constant multiplication, so it is not considered when identifying the previous layer.
If the previous layer is either a convolutional layer or an approximate ReLU layer, it may contain some zero values between each set of non-zero values. We can exploit this to efficiently execute the flatten operation by stacking the non-zero values using rotation and addition.
@@ IV-E2 Removing Invalid Data and Row Interval
If the previous layer is an average pooling layer, invalid values exist in the intervals between each set of valid data points. Moreover, the required constant multiplication (assuming a factor of $1/c^{2}$) has not been applied during the average pooling operation. Consequently, we must eliminate these invalid data points and perform the necessary multiplication. To address this issue, we suggest a three-step process as follows.
@@ IV-E4 Packing Multiple Ciphertexts
IV-E5 Time Complexity
The number of plaintext multiplications in the flatten layer of UniHENN is $\mathcal{O}(W_{in}+H_{in})$ and the number of rotations in the flatten layer of UniHENN is $\mathcal{O}(CH_{in}\cdot S_{total}+W_{in}+H_{in})$.
Therefore the complexity of flatten layer of UniHENN is $\mathcal{O}((CH_{in}\cdot S_{total}+W_{in}+H_{in})\cdot N\cdot\log N\cdot L^{2})$
@@ V-A Convolutional Layer
Let $W_{img}\times H_{img}$ be the image size and $N_{layer}$ the number of total convolutional layers. Denote $(W_{in(i)},H_{in(i)})$ and $(W_{out(i)},H_{out(i)})$ as the input and output sizes of the $i$-th convolutional layer, and the kernel size is $(W_{ker(i)},H_{ker(i)})$, stride size is $(W_{st(i)},H_{st(i)})$, and padding size is $(W_{pad(i)},H_{pad(i)})$ for all $i\in\left[1,N_{layer}\right]$. Additionally, let
$H_{[st(i)]}=\prod\limits_{j=1}^{i}H_{st(j)}$ (2)
Then, $H_{img}\geq H_{out(i)}\times H_{[st(i)]}$, where $H_{ker(i)}\geq H_{st(i)}\geq 1$ for all $i\in\left[1,N_{layer}\right]$.
@@
$H_{out(i)}=\left[\frac{H_{in(i)}-H_{ker(i)}}{H_{st(i)}}\right]+1$ (3)
@@
$H_{out(i)}\times H_{st(i)}$ $=\left(\left[\frac{H_{in(i)}-H_{ker(i)}}{H_{st(i)}}\right]+1\right)\times H_{st(i)}$ (4)
$=\left[\frac{H_{in(i)}-H_{ker(i)}+H_{st(i)}}{H_{st(i)}}\right]\times H_{st(i)}$ (5)
$\leq\left[\frac{H_{in(i)}}{H_{st(i)}}\right]\times H_{st(i)}\leq H_{in(i)}=H_{out(i-1)}$ (6)
@@
Note that after passing through the $i$-th convolutional layer, the interval in each column of data is $W_{img}\times(H_{[st(i)]}-1)$ (detailed in  IV-B). Given $W_{img}$ columns of data or fewer, the image size of $H_{img}\times W_{img}$ is not exceeded by Theorem 1.
@@
$W_{[st(i)]}=\prod\limits_{j=1}^{i}W_{st(j)}$ (7)
Then, $W_{img}\geq W_{out(i)}\times W_{[st(i)]}$, where $W_{ker(i)}\geq W_{st(i)}\geq 1$ for all $i\in\left[1,N_{layer}\right]$.
@@ V-B Average Pooling Layer
Assuming that the kernel size in the average pooling layer is $c$, each channel uses a filter with a kernel size of $(c,c)$ and a stride of $(c,c)$. The result is then multiplied by $1/c^{2}$. We omit the multiplication of a vector consisting of 0s and 1s, allowing some non-zero but unused values (i.e., invalid values) to overlap with other values. According to Theorem 1, the number of slots in the ciphertext data remains unchanged when passing through UniHENN’s average pooling. However, to prevent invalid values from occupying the positions of other used data, an additional space of $(W_{img}+1)\times(c-1)$ is required (detailed in IV-C).
@@ VI-A Experimental Settings
We used a server on the NAVER Cloud platform [44] with the following specifications: 16 vCPU cores (Intel(R) Xeon(R) Gold 5220 CPU @ 2.20GHz), 64GB of memory, and a 50GB SSD.
@@ VI-B Datasets
We utilized four different datasets for our models: MNIST [16], CIFAR-10 [29], USPS [21], and ECG [39].
@@ VI-B1 MNIST Dataset
The MNIST dataset [16] consists of 70,000 grayscale images of handwritten digits, each sized $28\times 28$ pixels, with 10 classes representing the digits from 0 to 9. We used 60,000 images for training and 10,000 images for testing.
@@ VI-B2 CIFAR-10 Dataset
The CIFAR-10 dataset [29] includes 60,000 RGB images, each sized $32\times 32$ pixels, with 10 classes. We used 50,000 images for training and 10,000 images for testing.